# The NEMESIS Catalogue of Young Stellar Objects for the Orion Star Formation Complex

### v1

@@ Appendix D Extragalactic and Giants contamination assessment
In order to probe the extragalactic and giant contamination in our catalogue, we built two binary Random Forest classifiers trained to recognise these types of sources based on a set of attributes and a labelled sample. Similar classifiers specialised at separating stellar and extragalactic sources have been extensively implemented in the literature (e.g., Vasconcellos et al. 2011; Logan & Fotopoulou 2020; Cook et al. 2024; Lourens et al. 2024). However, these previous studies do not typically include YSOs in their training sample. Hence here we build our own classifier tailored for the OSFC.
D.1 Data
@@
We built a reference photometric catalogue for the OSFC field based on large photometric surveys summarised in Sec. 2.2, with a base list of 16,177,367 sources with data in Gaia DR3, PanSTARRS1-DR2 and CatWISE. We have refrained from including near-infrared data because of the unavailability of a survey that homogeneously covered the full field of the OSFC with depth comparable with the other three surveys used. We refer to the RF classifier trained to recognize extragalactic sources against stellar sources as Extragalactic classifier, and the one trained to recognize giant stars against stellar and extragalactic sources, Giant classifier.
D.2 Labels
D.2.1 Known extragalactic sources
@@
We collated a sample of previously identified extragalactic sources, including a mixture of resolved and unresolved Galaxies, AGNs and QSOs. This accounts for 32,850 sources in the OSFC field. To maximise purity, we kept only extragalactic sources with a match within 0.5”-1” to the photometric reference catalogue. The list of extragalactic sources was composed by:
HyperLEDA database: 2,638 sources (Makarov et al. 2014);
Million Quasars catalogue v. 8: 4,641 sources (Flesch 2023);
Galaxy Zoo: 6,124 sources(Walmsley et al. 2023);
MANGROVE Catalogue: 1,396 sources (Biteau 2021)
Gaia DR3 high-purity Galaxy sample: 7,718 sources;
Gaia DR3 high-purity QSO sample: 15,629 sources;
@@
The extragalactic sources from Gaia DR3 were identified by Gaia’s ML classification and redshift estimation (Delchambre et al. 2023), brightness profile analysis (Ducourant et al. 2023), or variability behaviour (Carnerero et al. 2023; Rimoldini et al. 2023). High-purity samples were devised based on selection procedures outlined by Gaia Collaboration et al. (2023b).
D.2.2 Stellar sources
@@
The following distinct samples of stellar sources were labelled:
@@
Certainly young: 3,698 sources from the NEMESIS YSO catalogue which have spectroscopically confirmation of youth based on significant Lithium absorption in their spectra (Sec. 3.4);
@@
At Orion’s distance: 59,146 Gaia DR3 sources with parallax uncertainty better than 0.05 mas, a parallax consistent with a location at the surroundings of the OSFC (300-500 pc);
@@
Giants: 13,336 sources with in either RAVE, LAMOST LRS or MRS (Sec. 2.3), and precision better than 0.1 dex;
@@
Other Stellar Sources: 72,336 sources with in either RAVE, LAMOST LRS or MRS (Sec. 2.3), and precision better than 0.1 dex.
D.3 Features
We have carried out tests with different versions of the classifiers including as features all possible colours within our reference photometric catalogue and parameters such as Gaia’s RUWE, parallax, and coordinates. The final results included here used 24 features for the extragalactic classifier, including every possible photometric colour within the combination of PanSTARRS 5 bands (, , , , and ) and CatWISE 2 bands ( and ), and 4 colours including Gaia DR3’s G band. These features were available for 7,015,563 sources in the OSFC field. For the giant classifier, although we have tested the inclusion of other surveys, the best results were obtained while focussing on the data available from Gaia DR3. We included 6 features: parallax, , , , parallax_error, and , which were available for 8,385,766 sources in the OSFC field.
D.4 Training and results
@@
We have further processed the labelled data to remove any overlapping labels. For example, sources labelled both Certainly Young and Giant were removed. After these adjustments, and after the pre-cleaning of the data used in the classifiers, parent samples with 23,217 and 94,248 sources labelled extragalactic and stellar were available for the Extragalactic classifier, and 12,988 and 133,966 sources labelled giant and others were available for the Giant classifier. For each classifier, we used an 80-20% separation between training and testing samples.
For the extragalactic classifier, we used samples with 12,000 labelled extragalactic sources and twice as many stellar sources. For that, we have randomly selected sources from their parent samples while respecting certain proportions between sources coming from different sub-samples. For example, only resolved galaxies from Galaxy Zoo remained available after the pre-cleaning of the data, and we made sure that these composed at least about a sixth of the training and test samples.
@@
For the Giant classifier, we used 6,000 giants against a sample of non-giants composed of 2,500 certainly young sources, 5,000 other stellar sources and 1,000 extragalactic sources. We deliberately avoid the sample at Orion’s distance as this may contain Giants without measurements available to us. At the stellar sample, we made sure the sample always contained relevant fractions of certainly young sources and giants.
We implemented ensemble-based RF classifiers, composed of 50 RF binary classifiers trained with 50 samples randomly picked from the parent distributions discussed above. Results presented here are ensemble averages, with the confusion matrices in Fig. 16 and performance report in Tab. 13.
@@
| Extragalactic sources | 0.994 | 0.986 | 0.990 |
| Stellar sources | 0.993 | 0.997 | 0.995 |
@@
| Giants | 0.994 | 0.993 | 0.993 |
| Other sources | 0.995 | 0.996 | 0.995 |
The trained classifiers were used to derive probabilities of sources in the field of the OSFC being extragalactic or a giant. The distributions of probabilities estimated are shown in Fig. 17. The results of these classifiers allowed us to identify 327,557 probable extragalactic sources and 115,414 probable giants in the field of the OSFC based on a threshold of 90%. The sources for which we could derive such probabilities were matched to the NEMESIS Catalogue of YSOs in the OSFC and were used in Sec. 5.2 for investigation of the contamination levels by such sources within the catalogue.

### v2

@@ Appendix E Extragalactic and Giants contamination assessment
To investigate extragalactic and giant contamination in our catalogue, we used Scikit-learn (Pedregosa et al. 2011) to design two binary Random Forest classifiers (Breiman 2001) trained to disaggregate these types of sources based on a set of attributes and a labelled sample. Similar classifiers specialised in the separation of stellar and extragalactic sources have been extensively implemented in the literature (e.g., Vasconcellos et al. 2011; Logan & Fotopoulou 2020; Cook et al. 2024; Lourens et al. 2024). However, these previous studies do not explicitly include YSOs in their training sample. Hence, here, we build our own classifier tailored for the OSFC field.
E.1 Data
@@
We built a reference photometric catalogue for the OSFC field based on large photometric surveys summarised in Sect. 2.2, with a base list of 16 177 367 sources with data from Gaia DR3, PanSTARRS1-DR2 and CatWISE. Near-infrared data were excluded because there were no surveys covering the entire OSFC field with comparable depth to the other surveys. We refer to the RF classifier trained to identify extragalactic sources against stellar sources as “Extragalactic classifier”, and the one trained to identify giant stars against stellar and extragalactic sources, as “Giant classifier”.
E.2 Labelled dataset
E.2.1 Extragalactic sources
@@
We collated a sample of previously identified extragalactic sources, including a mixture of resolved and unresolved Galaxies, AGNs and QSOs. This accounts for 32 850 sources in the OSFC field. To maximise purity, we kept only extragalactic sources with a match within to the photometric reference catalogue. The list of extragalactic sources was composed of sources from:
HyperLEDA database (Makarov et al. 2014): 2 638 sources;
Million Quasars catalogue v. 8 (Flesch 2023): 4 641 sources;
Galaxy Zoo (Walmsley et al. 2023): 6 124 sources;
MANGROVE Catalogue (Biteau 2021): 1 396 sources;
Gaia DR3 high-purity Galaxy sample: 7 718 sources;
Gaia DR3 high-purity QSO sample: 15 629 sources.
@@
Extragalactic sources from Gaia DR3 were identified by Gaia’s ML classification and redshift estimation (Delchambre et al. 2023), brightness profile analysis (Ducourant et al. 2023), or variability behaviour (Carnerero et al. 2023; Rimoldini et al. 2023). High-purity samples were devised based on selection procedures outlined by Gaia Collaboration et al. (2023b).
E.2.2 Stellar sources
@@
The following samples of stellar sources were included in our labelled sample:
@@
Certainly young: 3 698 sources from the NEMESIS YSO catalogue which have spectroscopic confirmation of youth based on significant Lithium absorption in their spectra (Sect. 3.4);
@@
At Orion’s distance: 59 146 Gaia DR3 sources with parallax consistent with a location around the location of the OSFC (300-500 pc) and parallax uncertainty better than 0.05 mas;
@@
Giants: 13 336 sources with in either RAVE, LAMOST LRS or MRS (Sect. 2.3), and precision better than 0.25 dex;
@@
Other Stellar Sources: 72 336 sources with in either RAVE, LAMOST LRS or MRS (Sect. 2.3), and precision better than 0.25 dex.
E.2.3 Final sample
@@
The labelled data were further processed to remove any overlapping labels. For example, sources labelled both Certainly Young and Giant were removed. Further pre-cleaning of the data included removal of sources with very large uncertainties.
Extragalactic classifier sample:
All sources listed in Sect. E.2.1 were initially labelled as “extragalactic” (positive label), and all sources in Sect. E.2.2 were labelled as “stellar” (negative label). Validation of labelled samples was carried out by comparing the locus occupied by the different samples in colour-colour diagrams. During this procedure, we verified a significant tail of “Gaia DR3 high-purity QSOs” sources towards the location of stellar sources, which did not exist in the distribution of other samples with large numbers of QSOs (e.g., Million Quasars Catalogue). Furthermore, later classifier implementation tests revealed that the exclusion of “Gaia DR3 high-purity QSO” labelled greatly improved the classifier recall. We therefore excluded labels from the “Gaia DR3 high-purity QSO” sample from our final implementation. Further evidence that this sample suffers from contamination by YSOs is discussed in Sect. 5.2.3. A final parent sample with 23 217 and 94 248 sources labelled as extragalactic and stellar was available for the Extragalactic classifier.
Giant classifier sample:
@@
Sources in the Giants sample in Sect. E.2.2 were labelled as such (positive label), and extragalactic sources in Sect. E.2.1 or in the “certainly young” and “other stellar” samples were labelled as “other” (negative label). We deliberately avoid the sample at Orion’s distance as this may contain Giants with unavailable measurements. A final parent sample with 12 988 and 133 966 sources labelled as giant and others was available for the Giant classifier.
The final labelled datasets were split into 30% training, 40% validation and 30% testing sets. We ensured samples contained a representative fraction of the different source types (e.g., resolved and unresolved galaxies) by stratifying them based on origin catalogues (Sect. E.2.1), maintaining similar source proportions from each catalogue. Similarly, each subtype of stellar sources in Sect. E.2.2 was also adequately represented.
E.3 Features
Implementation tests were carried out by training classifiers in six subsets of features, combining available colours from Gaia DR3, PanSTARRS1-DR2, and CatWISE. We included both pure-colour subsets and versions with Gaia parallaxes and uncertainties. A summary of features used in each of our best-performing models is presented in Tab. .
| Model | Classifier | # features | features | sources |
[…]
| Gaia | Giant | 5 | , , , parallax, parallax_error | 8 385 766 |
E.4 Training
Class imbalance was intentionally maintained for both classifiers, as we expect giants and extragalactic sources to be a minority of sources in the OSFC field. However, this choice led to true negative labels dominating precision and accuracy. To reduce its impact on the smaller classes (true positive labels), we prioritised the use of scores that account for the trade-off between precision and recall for model comparison and performance evaluation. Additionally, as a general rule, models with validation sample true positive recovery rate under 75% were discarded. Grid search with 5-fold cross-validation was performed to optimise hyperparameters, using average precision as the reference score. Each classifier was then trained with the optimal parameters and exported.
E.5 Validation
Validation samples were used for the comparison of the performance of each classifier trained, and a series of scores (namely, accuracy, precision, recall, F1-score and F) score) was used to rank their performance. For the extragalactic classifier, we found that sets of attributes covering colours from optical to mid-infrared (GaiaPS1WISE_pure_colour) performed better, qualitatively consistent with results in previously published classifiers. For the giant classifier, the models using Gaia parallax performed better, with the Gaia model (attributes: Gaia colours plus parallax and its uncertainties) being only marginally outperformed by the GaiaWISE model (same Gaia attributes with the addition of CatWISE-based colours). Nevertheless, the requirement for both Gaia and CatWISE data for the latter makes it applicable to a smaller number of sources. We therefore opt to keep the results for the pure-Gaia model (Tab. ).
Validation samples were also used to study the precision-recall curves of the best-performing classifiers. At this stage, we used the F1-score for determining the best classification thresholds while ensuring a good trade-off between precision and recall. Best classification thresholds were found at 0.41 for the extragalactic classifier, and at 0.39 for the giant classifier.
E.6 Classification
We used the classification thresholds from the previous section for classification. Classification reports for the testing set are presented in Tab. 17, and confusion matrices in Fig. 17.
@@ E.6 Classification
| stellar | 1.00 | 1.00 | 1.00 |
| extragalactic | 0.99 | 0.97 | 0.98 |
@@
| Other | 0.99 | 0.99 | 0.99 |
| Giants | 0.94 | 0.93 | 0.93 |
Extragalactic classifier:
We were able to apply the GaiaPS1WISE_pure_colour model to classify 6 520 467 sources, with 139 668 classified as probable extragalactic sources. We note that among stellar sources misclassified as extragalactic (0.19 false negative rate), were spectroscopically confirmed YSOs, although these correspond to only of the YSOs in the testing sample.
Giant Classifier:
We were able to apply the Gaia model to derive the probabilities of being a giant for 8 385 766 sources in the field of the OSFC, with 141 808 likely giants identified. Among sources misclassified as giants in the testing sample, 97% were labelled “other stellar sources“ based on their measurements from spectroscopic surveys. The remaining 3% were real YSOs with spectroscopic constraints, these correspond to of the YSO sources in the testing sample.
The development of these classifiers included 53 154 labelled sources without enough data for classification. Both samples of classified and labelled but not classified sources were matched to the OSFC YSO catalogue and further discussed in Sect. 5.2.